# Interactive Learning of Physical Object Properties Through Robot Manipulation and Database of Object Measurements

### v1

@@ Abstract
Figure 1: Exploratory action selection to estimate object properties – schematic overview. Using the current belief, next action is selected to maximize the expected Information Gain (IG), taking the uncertainty of each action into account. After each action, new belief is computed using inference of Bayesian Network, and new action is selected.
@@ I Introduction
While rapid progress is being made in automatically extracting information about objects from large image datasets and text corpora, estimating physical properties remains a challenge. This work paves the way to automatic extraction of physical object properties such as material composition, mass, volume or stiffness, which typically require physically manipulating the object. We present a theoretical framework and its experimental evaluation that deals with exploratory action selection to maximize what can be learned about objects on the table in front of a robot manipulator. In our demonstrator, there is a partially adversarial mix of 17 household objects of different shapes and from different materials (incl. a real and a plastic banana, wine glass from glass and plastic, a soft and a hard sponge with identical appearance), 6 object properties (2 discrete – category and material; 4 continuous – elasticity, density, volume, and mass), and 5 actions the robot system can perform (2 based on vision – estimate the category or material from the image; 3 manipulation actions – squeeze, poke, and lift). We expressed the conditional dependencies between the properties using a Bayesian Network (BN) (see Fig. 2 and added freely available prior probability distributions as well as uncertainty associated with different measurement actions. After specifying which object property should be estimated, our algorithm chooses the best exploratory action based on expected IG, updates all object properties using inference in the BN, and picks the next optimal action to be executed. We demonstrate that the overall average performance of the system outperforms a baseline and analyze interesting runs showing how the algorithm automatically and efficiently discovers object properties. The robot pipeline is integrated with a logging module and an online database of objects, currently containing more than 24000 measurements of 63 different objects using four different grippers. We make publicly available all code and data, which provides a starting point for automatic digitization of objects and their physical properties by exploratory manipulations.
@@
Task and motion planning for robot manipulation, including sequences of actions, has been extensively studied using symbolic and recently also neuro-symbolic or neural approaches. In this work, the goal is not to perform a task (e.g., stack objects) but to learn about the properties of objects. Embodied reasoning for discovering object properties via manipulation was used in [1]; correct execution of a verbally specified task (e.g., stack objects by weight) required specific actions to learn about physical object properties. In this work, the order of actions is chosen not as to produce a cumulative result in the environment, but so that the action chosen next is optimal with respect to what can be learned about the object in front of the robot.
@@
Active manipulation for perception can be non-prehensile (e.g. “push to know” [2, 3]) or prehensile, i.e. grasping the object or exploring it with robot fingers (e.g., [4]). In most cases, manipulation serves object localization, shape reconstruction, or recognition; learning physical object properties has been less studied [5]. Haptic exploration of physical object properties using several actions and sensory modalities is extremely rare ([6] a notable exception). The exploratory movements are typically chosen based on some form of probabilistic model and information theoretic measures like entropy or IG (see e.g., [7]).
@@
An accompanying video is at https://youtu.be/h_ZIYUmzv-8; the code used in this work is available at https://github.com/ctu-vras/actsel; The object database and source code are available at https://cmp.felk.cvut.cz/ipalm/.
@@ II Exploratory Action Selection Driven by Information Gain
[…]
@@ II-A Initialization
Figure 3: Initialization of GMM from the reference data by weighting by corresponding probability mass function.
@@
Volume
[…]
Bottle Bowl Box Can Dice
$30\pm 50$
$40\pm 80$
$10\pm 26$
$8\pm 60$
$37\pm 50$
@@ II-B Inference
Following the execution of an action, the network is updated based on the obtained results (measurement update). These data points are then incorporated as a product into the distribution of the given property. For categorical properties, the posterior distribution post-update is propagated to all other nodes in the BN, adhering to the network’s connections as an undirected graph. This updating process, known as message passing, involves transmitting message vectors between nodes. The format and order of these message vectors correspond to the relevant probability functions: for interactions between nodes related to Material, the format aligns with the material PMF, while for other interactions, it aligns with the category PMF.
When the message originates from a continuous property node, such as Elasticity, the continuum must be first transformed into a finite vector to align with the required message format. Given our knowledge of the original reference distribution of materials for Density, we only need to estimate the weights of the GMM using the Expectation–maximization (EM) algorithm. This approach allows the classification of unknown objects into material categories without necessitating a class for every known material. Consequently, objects that fall between two classes can be represented, for instance, as a blend, like $0.5\times property_{1},0.5\times property_{2}$. Subsequently, this message vector from Elasticity serves as a virtual measurement for the Material node. Posterior probabilities are computed using a confusion matrix to simulate the uncertainty inherent in message passing. Specifically, this confusion matrix is unique to each connection between nodes; it is computed beforehand at the initialization step from the reference data for every blue arrow in the network model depicted in Fig. 2.
@@ II-C Measurement Models
The action selection algorithm needs to estimate the consequences of performing different actions, i.e. measurements. A measurement model, i.e. the accuracy of every measurement is required. For categorical variables, confusion matrices are used to compute the expected measurement. For brevity, only the average accuracy for each categorical action is given. Complete confusion matrices are available in the code repository. For continuous variables, a measurement is modeled as a normal distribution. The measurement accuracies and SDs are provided in the Table III. The SD for density was experimentally obtained based on the average performance of extracting a correct density belief from a mass measurement and a volume belief on the available experimental data. The mass measurement with our experimental setup has an accuracy of $\pm$ 6 [grams] for up to a 1 [kg] payload.
@@ II-D Expected Information Gain
In our framework, IG quantifies the reduction in entropy of the probability distribution of one or more variables in our BN after executing an action. It reflects how much information the action and its associated measurement have provided. When we compute the difference in entropy before and after conducting an action, we refer to it as experimental or true IG. While experimental IG offers insight into the performance of an algorithm, by definition, it alone cannot be used to decide which action to take. Therefore, to guide the decision-making, one must create an emulation—an estimation of how the property distribution will likely appear after each action—to choose the action most likely to reduce entropy, thus increasing IG. The experimental IG is formalized as:
@@
where $\nu$ represents a distribution of a node, either continuous or categorical (e.g., Elasticity: PDF or Category: PMF), $\Delta H(\nu)$ denotes the experimental IG, which is the difference in entropies. Here, $H(\nu{{}_{k}})$ is the entropy of a node distribution at algorithm iteration $k$, representing the prior, and $H(\nu_{k+1})$ is the entropy of the same node distribution at the next iteration step $k+1$, representing the posterior after the measurement. The expected IG is formalized as follows
@@
where $\mathbb{E}[\Delta H(\nu)]$ is the expected IG, and $\mathbb{E}[H(\bar{\nu_{k}})]$ is the expected entropy of the posterior distribution, which is the entropy of the emulation. Note that the emulation also contains the index $k$, not $k+1$, to emphasize that it reflects only an expectation of the posterior based on the current belief. For the continuous properties, theoretically, the emulation could involve sampling infinitely many possible scenarios of the action measurements based on the probability density function of a given property. However, such an approach would require simulating Bayesian updates for a large number of samples, which consumes considerable CPU time. Since the variance of the action related to a given property is available, we chose a more efficient option—convolving the continuous prior probability density function of a given property $\nu_{k}$ with a normal distribution having a zero mean and variance equal to the variance of the action measurement in question $\mathcal{N}(0,\sigma_{a}^{2})$. This procedure propagates the expected error across the entire continuous probability density function of the property. Formally:
@@
Subsequently, the expected continuous entropy is numerically calculated as:
@@
where the entropy is computed for each step on the support $\mathcal{X}$ of the property distribution. $h_{\mathcal{X}}$ denotes the normalization factor, determined by the bin size used for computing the sum-integral approximation. Strictly mathematically speaking, the Eq. (4) provides an integral approximation.
@@ 9             $\mathbb{E}_{k}[\Delta H(N)]:=\eta_{k}-\overline{\eta_{k}}$
Algorithm 1 Action selection (ACSEL)
@@ IV Experiments and Results
In our experiments, one of 17 objects was placed in front of the robot. There were 5 actions or measurements (cat-vision, mat-vision, mat-sound, weighing, squeezing) in the repertoire that could be selected by our algorithm based on their expected IG, i.e. their likelihood to maximally reduce the uncertainty in one or more of the object properties. There were five object properties (2 discrete variables with PMFCategory and Material; 3 continuous with PDFElasticity, Density, Volume/Mass)—see Fig. 2 for an overview. The action/measurement informing the Category node was cat-vision; mat-vision or mat-sound could be employed to update the Material node; squeezing measured Elasticity. Density and Volume nodes did not have a corresponding measurement that would update them directly and relied on relationships within the BN.
@@
Optimization modes. To drive the action selection, one of the nodes could be selected for which the uncertainty should be reduced (say the user is interested in a specific object property). For Category or Material, the expected discrete IG for every action was be evaluated and the best action executed. For the continuous variables (Elasticity, Density, Volume), the corresponding differential IGs were evaluated. We have also assessed runs optimizing Category and Material together or the three continuous property nodes together. Aiming to reduce uncertainty in all five nodes at the same time would require adding the discrete and differential IG which is not mathematically rigorous, so this regime was not evaluated. Please note that all actions have their associated measurement models, i.e. the system has an estimate of the noise associated with every action. Thus, optimizing one property does not always lead to executing the action that directly measures that property.
@@ IV-A Qualitative Illustration of the System’s Operation
The operation of the system is illustrated in Fig. 5 on three selected objects. In the first row (Panels A1-A4), the YCB wineglass, which is plastic, is explored while the Category variable is optimized. Cat-vision has the highest expected IG (panel A2) and is always chosen (panel A3). In the single run (panel A2), additional steps of the algorithm are shown. However, in this particular run, if termination was on, no more action would be executed as all actions have negative expected IG at step 1. Panel A4, right, shows the correct identification of the object Category as wineglass.
The other two rows show exploration of a sponge—a hardened exemplar in the middle (B1-B4) and a soft one at the bottom (C1-C4) (see accompanying video at https://youtu.be/h_ZIYUmzv-8);, in both cases with learning about Elasticity as the objective. Although the goal was to learn about the object stiffness, Cat-vision was chosen as the first most informative action to serve this objective. The squeezing action executed in the next step (see B2 and B3) adds information about the object being stiff, as is apparent in the final elasticity distribution as well as the material distribution (B4 vs. C4).
Figure 5: System operation and performance – examples. (Left – A1/B1/C1) Object name and image. (A2/B2/C2) Single run of experiment with termination criterion off. Expected IG of the actions shown. Triangle markers show which action was chosen in a given step. The dashed black line shows the experimental true IG after the action was executed. This action is not shown anymore in subsequent algorithm steps. (A3/B3/C3) Count of actions chosen during the runs with the termination on (if all available actions at a given step had negative expected IG, execution stopped). (A4/B4/C4) Probability distributions of relevant variables of the single run shown in (A2/B2/C2).
@@ IV-B Action Statistics and Intelligent Exploration Termination
Fig. 6 shows how many times different actions were executed when different object properties, i.e. different nodes in the BN, were selected to be “learned about”. When reducing the uncertainty in the Category node is the target, Panel A shows that cat-vision was always selected as the first action. In about half the cases, the algorithm then terminated. In the other cases, mat-vision was picked in the second step. When asked to learn about the Material (Panel C), interestingly, the algorithm picked the Weighing action as the most effective in all cases, rather than the mat-vision or mat-sound actions that directly target the Material node through their measurement update. This illustrates how the algorithm automatically takes the complete BN and the measurement uncertainties into account. For the continuous variables (Panels B, D), the algorithm typically performed all five actions. Cat-vision wass the most informative action to learn about Category (Panel A), Elasticity (B), and all the continuous together variables (D) but not about Material (C).
@@ IV-C Average performance across all objects
To demonstrate the effectiveness of the proposed algorithm over all the objects, we ran the algorithm with 5 repetitions on each of the 17 objects, using the action selection algorithm proposed optimizing the Category node and with random action selection as a baseline (i.e., 2 action selection regimes x 17 objects x 5 repetitions x 5 algorithm steps). For comparison purposes, the termination check was off. Fig. 7 shows the average entropy of the Category variable after 5 steps of the algorithm in the two action selection regimes, including the variance. The intelligent action selection (ACSEL) outperforms RAND. Furthermore, in the ACSEL regime, the plot shows a dramatic drop in entropy after the execution of the first action/measurement, demonstrating that the most effective action was chosen in the first step for the majority of the objects.
Figure 7: Average entropy (lower is better) of the Category node – Optimizing Category, 17 objects, 5 repetitions each. Comparing action selection based on IG (ACSEL) and random action selection (RAND).
To complement the analysis in Fig. 8 that shows entropy, which can be measured autonomously by the system, Fig. 8 displays cross-entropy which considers how well the current probability distribution of a node corresponds to the ground truth—in this case the true label for a given Category or Material. For both properties, ACSEL on average outperforms RAND. The first algorithm step is on average effective in reducing the cross-entropy. Additional measurements may, in fact, add noise to the system, increasing the entropy of the PDFs. However, in these cases, these steps will often not be executed (see Fig. 6 A,C).
Figure 8: Cross-entropy of the Category and Material nodes – Optimizing Category, 17 objects, 5 repetitions each. Comparing action selection based on IG (ACSEL) and random action selection (RAND).

### v2

@@ Abstract
Figure 1: Exploratory action selection to estimate object properties – schematic overview. Using the current belief, next action is selected to maximize the expected Information Gain (IG), taking the uncertainty of each action into account. After each action, new belief is computed using inference in the Bayesian Network, and a new action is selected.
@@ I Introduction
While rapid progress is being made in automatically extracting information about objects from large image datasets and text corpora, estimating physical properties remains a challenge. This work paves the way to automatic extraction of physical object properties such as material composition, mass, volume or stiffness, which typically require physically manipulating the object. We present a theoretical framework and its experimental evaluation that deals with exploratory action selection to maximize what can be learned about objects on the table in front of a robot manipulator (see Fig. 1 for an overview). In our demonstrator, there is a partially adversarial mix of 17 household objects of different shapes and from different materials (incl. a real and a plastic banana, wine glass from glass and plastic, a soft and a hard sponge with identical appearance), 6 object properties (2 discrete – category and material; 4 continuous – elasticity, density, volume, and mass), and 5 actions the robot system can perform (2 based on vision – estimate the category or material from the image; 3 manipulation actions – squeeze, poke, and lift). We expressed the conditional dependencies between the properties using a Bayesian Network (BN) (see Fig. 2) and added freely available prior probability distributions as well as the uncertainty associated with different measurement actions. After specifying which object property should be estimated, our algorithm chooses the best exploratory action based on expected Information Gain (IG), updates all object properties using inference in the BN, and picks the next optimal action to be executed. We demonstrate that the overall average performance of the system outperforms a baseline and analyze interesting runs showing how the algorithm automatically and efficiently discovers object properties. The robot pipeline is integrated with a logging module and an online database of objects, currently containing more than 24000 measurements of 63 different objects using four different grippers. We make all code and data publicly available, which provides a starting point for automatic digitization of objects and their physical properties by exploratory manipulations.
@@
Task and motion planning for robot manipulation, including sequences of actions, has been extensively studied using symbolic and recently also neuro-symbolic or neural approaches. In this work, the goal is not to perform a task (e.g., stack objects) but to learn about the properties of objects. Embodied reasoning for discovering object properties via manipulation was used in [1]; correct execution of a verbally specified task (e.g., stack objects by weight) required specific actions to learn about physical object properties. In this work, the order of actions is chosen not to produce a cumulative result in the environment, but so that the action chosen next is optimal with respect to what can be learned about the object in front of the robot.
@@
Active manipulation for perception can be non-prehensile (e.g. “push to know” [2, 3]) or prehensile, i.e. grasping the object or exploring it with robot fingers (e.g., [4]). In most cases, manipulation serves object pose estimation (e.g., [5]), shape reconstruction, or recognition; learning physical object properties has been less studied (see [6, 7] for reviews). Haptic exploration of physical object properties using several actions and sensory modalities is extremely rare (with [8] a notable exception). The exploratory movements are typically chosen based on some form of probabilistic model and information theoretic measures like entropy or IG (see e.g., [9, 10]).
@@
An accompanying video is available at https://youtu.be/h_ZIYUmzv-8; the code used in this work is hosted at https://github.com/ctu-vras/actsel; The database of object measurements and its source code can be found at https://cmp.felk.cvut.cz/ipalm/.
@@ II Exploratory Action Selection Driven by Information Gain
Elasticity is a continuous variable expressed in Young’s modulus [kPa], Density in homogeneous density [kg/$m^{3}$], and Volume in the volume of the object’s material in [m3].
@@ II-A Initialization
Figure 3: Initialization of GMM from the reference data – weighting by corresponding probability mass function.
@@
Volume [cm3] Bottle Bowl Box Can Dice
$30\pm 50$ $40\pm 80$ $10\pm 26$ $8\pm 60$ $37\pm 50$
@@ II-B Inference
Following the execution of an action, the network is updated based on the obtained results (measurement update). These data points are then incorporated as a product into the distribution of the given property. For categorical properties, the posterior distribution update is propagated to all other nodes in the BN, adhering to the network’s connections as an undirected graph. This updating process, known as message passing, involves transmitting message vectors between nodes. The format and order of these message vectors correspond to the relevant probability functions: for interactions between nodes related to Material, the format aligns with the material PMF, while for other interactions, it aligns with the category PMF.
When the message originates from a continuous property node, such as Elasticity, the continuum must be first transformed into a finite vector to align with the required message format. Given our knowledge of the original reference distribution of materials for Density, we only need to estimate the weights of the GMM using the Expectation–maximization (EM) algorithm. This approach allows the classification of unknown objects into material categories without necessitating a class for every material. Consequently, objects that fall between two classes can be represented, for instance, as a mixture, like $0.5\times property_{1},0.5\times property_{2}$. Subsequently, this message vector from Elasticity serves as a virtual measurement for the Material node. Posterior probabilities are computed using a confusion matrix to model the inherent uncertainty. Specifically, this confusion matrix is unique to each connection between nodes; it is determined beforehand at the initialization step from the reference data for every blue arrow in the network model depicted in Fig. 2.
@@ II-C Measurement Models
The action selection algorithm needs to estimate the consequences of performing different actions, i.e. measurements. A measurement model specifying the accuracy of every measurement is required. For categorical variables, confusion matrices are used to compute the expected measurement. For brevity, only the average accuracy for each categorical action is given. Complete confusion matrices are available in the code repository. For continuous variables, a measurement is modeled as a normal distribution. The measurement accuracies and SDs are provided in the Table III. The SD for density was experimentally obtained based on the average performance of extracting a correct density belief from a mass measurement and a volume belief on the available experimental data. The mass measurement with our experimental setup has an accuracy of $\pm$ 6 [grams] for up to a 1 [kg] payload.
@@ II-D Expected Information Gain
In our framework, IG quantifies the reduction in entropy of the probability distribution of one or more variables in our BN after executing an action. It reflects how much information the action and its associated measurement have provided. When we compute the difference in entropy before and after conducting an action, we refer to it as experimental or true IG. While experimental IG offers insight into the performance of an algorithm, by definition, it alone cannot be used to decide which action to take. Therefore, to guide the decision-making, one must create an emulation—an estimation of the property distribution after each action—to choose the action most likely to reduce entropy, thus increasing IG. The experimental IG is formalized as:
@@
where $\nu$ represents a distribution of a node, either continuous or categorical (e.g., Elasticity: PDF or Category: PMF), $\Delta H(\nu)$ denotes the experimental IG, which is the difference in entropies. Here, $H(\nu{{}_{k}})$ is the entropy of a set of nodes at algorithm iteration $k$, representing the prior, and $H(\nu_{k+1})$ is the entropy of the same set of nodes at the next iteration step $k+1$, representing the posterior after the measurement. The expected IG is formalized as follows:
@@
where $\mathbb{E}[\Delta H(\nu)]$ is the expected IG, and $\mathbb{E}[H(\bar{\nu_{k}})]$ is the expected entropy of the posterior distribution, which is the entropy of the emulation. Note that the emulation also contains the index $k$, not $k+1$, to emphasize that it reflects only an expectation of the posterior based on the current belief. Theoretically, the emulation for continuous properties can be achieved through sampling. However, this is computationally prohibitively expensive. Since the variance of the action related to a given property is available, we chose a more efficient option—convolving the continuous prior probability density function of a given property $\nu_{k}$ with a normal distribution having a zero mean and variance equal to the variance of the action measurement in question $\mathcal{N}(0,\sigma_{a}^{2})$. This procedure propagates the expected error across the entire continuous probability density function of the property. Formally:
@@
Subsequently, the expected continuous entropy is numerically approximated as:
@@
where the entropy is computed for each step on the support $\mathcal{X}$ of the property distribution. $h_{\mathcal{X}}$ denotes the normalization factor, determined by the bin size used for computing the sum-integral approximation. Strictly mathematically speaking, Eq. (4) provides an integral approximation.
@@ 13      $a_{k}^{*}\in\textit{argmax}_{a\in\mathcal{A}_{k}}\mathbb{E}_{k}[\Delta H(N)]$
Algorithm 1 Action selection (ACTSEL)
@@ IV Experiments and Results
In our experiments, one of 17 objects was placed in front of the robot. There were 5 actions or measurements (cat-vision, mat-vision, mat-sound, weighing, squeezing) in the repertoire that could be selected by our algorithm based on their expected IG, i.e. their likelihood to maximally reduce the uncertainty in one or more of the object properties. There were five object properties (2 discrete variables with PMFCategory and Material; 3 continuous with PDFElasticity, Density, Volume/Mass)—see Fig. 2 for an overview. The cat-vision was the action/measurement informing the Category node; mat-vision or mat-sound could be employed to update the Material node; squeezing measured Elasticity. Density and Volume nodes did not have a corresponding measurement that would update them directly and relied on relationships within the BN.
@@
Optimization modes. To drive the action selection, one of the nodes could be selected for which the uncertainty should be reduced (say the user is interested in a specific object property). For Category or Material, the expected discrete IG for every action was evaluated to select the best action. For the continuous variables (Elasticity, Density, Volume), the corresponding differential IGs were evaluated. We have also assessed runs optimizing Category and Material together or the three continuous property nodes together. Aiming to reduce uncertainty in all five nodes at the same time would require calculating a weighted sum of the discrete and differential IG, which is not mathematically rigorous, so this regime was not evaluated. Please note that all actions have their associated measurement models, i.e. the system has an estimate of the noise associated with every action. Thus, optimizing one property does not always lead to executing the action that directly measures that property.
@@ IV-A Qualitative Illustration of the System’s Operation
The operation of the system is illustrated in Fig. 5 on three selected objects. In the first row (Panels A1-A4), the YCB wineglass, which is plastic, is explored while the Category variable is optimized. Cat-vision has the highest expected IG (panel A2) and is always chosen first (panel A3). In the single run (panel A2), additional steps of the algorithm are shown. However, in this particular run, if termination was on, no more action would be executed as all actions have negative expected IG at step 1. Panel A4, right, shows the correct identification of the object Category as wineglass.
The other two rows show exploration of a sponge—a hardened exemplar in the middle (B1-B4) and a soft one at the bottom (C1-C4) (see accompanying video at https://youtu.be/h_ZIYUmzv-8); in both cases with learning about Elasticity as the objective. Although the goal was to learn about the object’s stiffness, Cat-vision was chosen as the first most informative action to serve this objective. The squeezing action executed in the next step (see B2 and B3) adds information about the object being stiff, as is apparent in the final elasticity distribution as well as the material distribution (B4 vs. C4).
Figure 5: System operation and performance – examples. (Left – A1/B1/C1) Object name and image. (A2/B2/C2) Single run of experiment with termination criterion off. Expected IG of the actions shown. Triangle markers show which action was chosen in a given step. The dashed black line shows the experimental true IG after the action was executed. This action is not shown anymore in subsequent algorithm steps. (A3/B3/C3) Count of actions chosen during the runs with the termination on (if all available actions at a given step had negative expected IG, execution stopped). (A4/B4/C4) Final probability distributions of relevant variables of the single run shown in (A2/B2/C2).
@@ IV-B Action Statistics and Intelligent Exploration Termination
Fig. 6 shows how many times different actions were executed when different object properties, i.e. different nodes in the BN, were selected to be “learned about”. When reducing the uncertainty in the Category node, cat-vision was always selected as the first action (Panel A). In about half the cases, the algorithm then terminated. In the other cases, mat-vision was picked in the second step. When asked to learn about the Material (Panel C), interestingly, the algorithm picked the Weighing action as the most effective in all cases, rather than the mat-vision or mat-sound actions that directly target the Material node through their measurement update. This illustrates how the algorithm automatically takes the complete BN and the measurement uncertainties into account. For the continuous variables (Panels B, D), the algorithm typically performed all five actions. Cat-vision wass the most informative action to learn about Category (Panel A), Elasticity (B), and all the continuous together variables (D) but not about Material (C).
@@ IV-C Average performance across all objects
To demonstrate the effectiveness of the proposed algorithm over all the objects, we ran the algorithm with 5 repetitions on each of the 17 objects, using the action selection algorithm proposed optimizing the Category node and with random action selection as a baseline (i.e., 2 action selection regimes x 17 objects x 5 repetitions x 5 algorithm steps). For comparison purposes, the termination check was off. Fig. 7 shows the average entropy of the Category variable after 5 steps of the algorithm in the two action selection regimes, including the variance. The intelligent action selection (ACTSEL) outperforms RAND. Furthermore, in the ACTSEL regime, the plot shows a dramatic drop in entropy after the execution of the first action/measurement, demonstrating that the most effective action was chosen in the first step for the majority of the objects.
Figure 7: Average entropy (lower is better) of the Category node – Optimizing Category, 17 objects, 5 repetitions each. Comparing action selection based on IG (ACTSEL) and random action selection (RAND).
To complement the analysis in Fig. 7 that shows entropy, which can be measured autonomously by the system, Fig. 8 displays cross-entropy which considers how well the current probability distribution of a node corresponds to the ground truth—in this case the true label for a given Category or Material. For both properties, ACTSEL on average outperforms RAND. The first algorithm step is on average effective in reducing the cross-entropy. Additional measurements may, in fact, add noise to the system, increasing the entropy of the PDFs. However, in these cases, these steps will often not be executed (see Fig. 6 A,C).
Figure 8: Cross-entropy (lower is better) of the Category and Material nodes, i.e. comparing estimation with ground truth (optimizing Category, 17 objects, 5 repetitions each). Comparing action selection based on IG (ACTSEL) and random action selection (RAND).